what is i(root over -1) to the power i?what type of complex no it will be?and pls tell me something about logarithm of complex numbers

Grade:12

2 Answers

Chetan Mandayam Nayakar
12 years ago

MY PREVIOUS SOLUTION WAS WRONG.

(√(-1))i=e(i(((pi/2)+2n*pi)))*i, where n is any integer. Therefore, there are infinitely many values.

jitender lakhanpal
12 years ago

dear Debanuj

i to the power i = e-∏/2 

where 

e= 2.7108   and    ∏ = 22/7

 proof 

let 

y = i to the power i 

taking natural logarithms on both sides (ln is log base e)

ln y = i ln i 

changing i into eulerian form

i = e to the power i∏/2

ln y = i*i∏/2

as i*i = -1

lny = -∏/2

y = e to the power -∏/2

so proved 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  

jitender 

askiitians expert